Mr. Jackson is 13 years less than 3 times as old as his son. Six years ago, he was 14 more than twice as old as his son then. Find each of their ages.

Answer by stanbon(75887)

Mr. Jackson is 13 years less than 3 times as old as his son. Six years ago, he was 14 more than twice as old as his son then. Find each of their ages.
----
Equations:
J = 3S - 13
J-6 = 2(S-6)+14
-----
Substitute for "J" and solve for "S.
3S-13-6 = 2S-12+14
----
S = -8+19
Son's age = 11 yrs
---
Solve for "J"::
J = 3S-13
J = 33-13 = 20 yrs (Mr. Jackson's age now)
----------------------
Cheers,
Stan H.
----------------------